A current I produces a magnetic field B at the centre of the circular coil. If the coil rewound to double its number of turns then what is the magnetic field at its centre in terms of B?
1
Expert's answer
2020-07-13T12:21:53-0400

Answer:

The magnetic effect of a current coil depends on the number of turns, and the stronger the more turns the coil has.

If you double the number of turns, the magnetic field B in the center of the circular coil, with constant current will double.

The magnetic induction inside the long solenoid is equal to


"{\displaystyle B={\frac {4\pi }{c}}{\frac {NI}{l}}};"


where B is the magnetic induction, c is the speed of light, N is the number of turns, l is the length of the coil.